# Implicit Geometric Descriptor-Enabled ANN Framework for a Unified Structure-Property Relationship in Architected Nanofibrous Materials

### v1

@@ 3.3 Generalization capability of implicit geometric descriptor-enabled ANN framework
To demonstrate that the unified structure property relationship we developed could serve as a predictive model to predict the mechanical properties of previously unknown architectures, we designed several new architectural designs—both cylinder-based and sinusoid-based—ensuring they were distinct from the 81 architectures used to train the ANN. However, we kept the descriptor values of these new architectures within the range of those 81 architectures used in the training of the ANN. The new cylinder-based designs were created by incorporating concentric rings within fractal arrays, resulting in four distinct architectural types ($C_{1}-C_{4}$)(Fig. 5(a-d)): $C_{1}$ and $C_{2}$ are first-order fractal arrays where $C_{1}$ has concentric rings applied to the inscribed cylinders while $C_{2}$ has concentric rings applied to both the inscribed and circumscribed cylinders. $C_{3}$ and $C_{4}$ are second-order fractal arrays where $C_{3}$ has concentric rings applied to the first-order inscribed cylinders while $C_{4}$ has concentric rings applied to both the first- and second-order inscribed cylinders as well as the circumscribed cylinder. The sinusoid-based designs, included 5 different types ($S_{1}-S_{5}$) (Fig. 5(e-i)): $S_{1}$ and $S_{2}$ have sinusoids running in single direction while $S_{3}$, $S_{4}$, and $S_{5}$ employ orthogonal sinusoid patterns with sinusoids running in both directions. $S_{2}$ has the same sinusoidal pattern as $S_{1}$ with additional cylinders in the gaps, and $S_{5}$ has additional cylinders within the gaps of $S_{3}$. $S_{4}$ differs from $S_{3}$, using a different sine function for its design.
The ANN models gave accurate predictions for all three specific mechanical properties for the architectures used to demonstrate it’s predictive capability. In Fig. 5(j-l), the predicted values of these mechanical properties are compared with experimental values, demonstrating the model’s accuracy. Fig. 5(m) summarizes this by presenting the ratio ($\eta$) of predicted to experimental values for $W^{*}$, $\sigma^{*}$, and $E^{*}$ across designs $C_{1}$ to $S_{5}$. Notably, the prediction performance for the sinusoid-based architectures was not only excellent but slightly better than that of most of the cylinder-based ones. This is a result of the descriptor values of sinusoid-based architectures being closer to the expected values of the distribution of the descriptors used during training. This outcome suggests that our ANN approach, based on implicit geometric descriptors, can be generalized to predict the specific mechanical properties of arbitrary architected VACNT foams as well.
@@ 3.4 Implicit geometric descriptor-based design for independent tuning of mechanical properties
To guide the implicit geometric descriptor based design of architected VACNT foams, as a first step, we quantified the individual influence of each descriptor on the mechanical properties, through a “quotient of error increase” sensitivity analysis [45] for our ANN model. This method measures how much the network’s error increases when a specific descriptor is excluded, compared to when all descriptors are used. It enables us to assess the importance of each descriptor by evaluating the impact of its absence on the model’s overall performance. For this analysis, we trained the model separately for each mechanical property of interest and computed the MSE and $R^{2}$ values using only one of the two geometric descriptors (Fig. 6(a)). The descriptor importance score $\Omega_{\mathcal{X}}$ is the ratio between the MSE obtained when the descriptor $\mathcal{X}$ is excluded—$\mathrm{MSE}_{\mathcal{X}}$—and the MSE for the original network (Equation 16), providing a measure of the model’s sensitivity to $\mathcal{X}$. Comparing $\Omega$ values across descriptors reveals their relative importance to the ANN model, as shown in Fig. 6(b).
@@
The consistently higher values obtained for $\Omega_{\mathcal{D}}$ across all three specific mechanical properties indicates that $\mathcal{D}$ is the more influential parameter of the two in determining the mechanical properties (Fig. 6(b)). This suggests that the compactness of the geometries in the architecture should be prioritized to achieve higher specific mechanical properties. However, notably, the influence of $\mathcal{D}$ is lower on $\sigma^{*}$ and even lesser on $E^{*}$, despite $\Omega_{\mathcal{I}_{mc-c}}$ remaining almost unchanged (Fig. 6(b)). This trend is also reflected in the $R^{2}$ values computed for the ANN when one of the geometric descriptors is excluded. The $R^{2}$ values are consistently higher in the absence of $\mathcal{I}_{mc-c}$, although they are reduced for $\sigma^{*}$ and $E^{*}$ (Fig. 6(c)). This finding indicates that $W^{*}$ is predominantly governed by the nanoscale interactions among the nanotubes—captured by $\mathcal{D}$—whereas there is a strong interplay between size effect—captured by $\mathcal{I}_{mc-c}$—and the nanotube interactions in governing the mechanical properties $\sigma^{*}$ and $E^{*}$. Our multi-variate regression analysis with interdependence term included (MVLR-I) further supports this statistically significant interdependence between $\mathcal{D}$ and $\mathcal{I}_{mc-c}$, with p-values of 0.00027 for $\sigma^{*}$ and 0.00015 for $E^{*}$. These findings imply that densely packed and aligned CNTs within confined regions enhance nanotube interactions upon loading, reinforcing the critical role of compactness on $\sigma^{*}$ and $E^{*}$. Therefore, optimizing both $\mathcal{I}_{mc-c}$ and $\mathcal{D}$ in the design of architected VACNT foams could enhance strength and modulus, while designs focused solely on $\mathcal{D}$ may maximize energy absorption. This sensitivity analysis revealed the significance of each descriptor on different mechanical properties, providing a potential foundation for designing VACNT foams to achieve various independent property combinations tailored for a wide range of applications.
@@ 3.5 Generalizable design strategies for target mechanical properties
Understanding how tuning $\tau$ and $\gamma_{i}$ maps onto the mechanical property landscape further refines our strategy. Although $\tau$ may influence $\mathcal{I}_{mc-c}$ differently depending on the region in the $\mathcal{D}$ space, it is intuitive that increasing $\tau$ generally reduces $\sigma^{*}$ and $E^{*}$ owing to the size effect and the reduction in the unit cell population as the unit cell size increases. On the other hand, increasing $\tau$ decreases $\mathcal{D}$ (Fig. 1(d))—the predominant influencer of $W^{*}$—resulting in enhancement of $W^{*}$ (Fig. 6(d)). These trends collectively suggest that by increasing $\tau$, periodic architectures with low $\sigma^{*}$ and $E^{*}$ can be achieved while maintaining high $W^{*}$. Fig. 7(c) and Fig. 7(d) respectively illustrate the $\sigma^{*}$ vs. $W^{*}$ and $E^{*}$ vs. $W^{*}$ property landscapes for cylindrical architectures, highlighting constant $t$ (analogous to $\tau$) contours as predicted by the ANN model. As $t$ increases, the contours shift towards right and stays below the smaller $t$ contours indicating higher $W^{*}$ and lower $\sigma^{*}$ and $E^{*}$. However, along each contour, all three mechanical properties simultaneously increase in the direction where $\mathcal{I}_{mc-c}$ and $\mathcal{D}$ simultaneously decrease due to variations in parameters such as $D_{in}$ and $g$ (analogous to $\gamma_{i}$). Guided by these behaviors, we propose a generalizable, qualitative $\sigma^{*}$(or $E^{*}$) vs. $W^{*}$ property landscape (Fig. 7(e)) for an arbitrary unit cell-based periodic architecture (Fig. 7(e)(Inset)). This landscape presents qualitative constant $\tau$ contours capturing the insights gained from the cylindrical architectures. With the selection of an appropriate $\tau$ to define a target property regime, the $\mathcal{I}_{mc-c}$ and $\mathcal{D}$ can be adjusted by fine-tuning $\gamma_{i}$ to navigate along the contours to realize desired combination of mechanical properties. By leveraging this implicit geometric descriptor enabled-design framework one can simplify the design process and realize systematic pathways for tailoring the mechanical performance of architected VACNT foams across a broad spectrum of applications.
@@ Acknowledgments
MATLAB codes for generating patterns and computing MCSIs are available on https://github.com/ThevamaranLab/MCSI-Computation
@@
Fig. S1 available in Supplementary information (PDF)
Table S1, S2 , S3 and S4 available in Supplemental information (PDF)
Data S1. Data relevant to ANN framework, architecture patterns and MCSIs (Compressed zip folder)

### v2

@@ 3.3 Generalization capability of implicit geometric descriptor-enabled ANN framework
[…]
The ANN models gave accurate predictions for all three specific mechanical properties for the architectures—$C_{1}$ to $S_{5}$—used to demonstrate it’s predictive capability. In Fig. 5(m-o), the predicted values of these mechanical properties are compared with experimental values, demonstrating the model’s accuracy. Fig. 5(p) summarizes this by presenting the ratio ($\eta$) of predicted to experimental values for $W^{*}$, $\sigma^{*}$, and $E^{*}$ across designs $C_{1}$ to $S_{5}$. Notably, the prediction performance for the sinusoid-based architectures was not only excellent but slightly better than that of most of the cylinder-based ones. This is a result of the descriptor values of sinusoid-based architectures being closer to the expected values of the distribution of the descriptors used during training. This outcome suggests that our ANN-enabled structure property relationship—based on implicit geometric descriptors—is generalizable across a wide spectrum of architectures.
The minor deviations in the predictions highlight the higher order effects emerging from the architectural intricacies—subtleties that extend beyond the first-order nanoscale mechanism driven trends captured by the two MCSI descriptors employed in this framework. To account for this, the flexibility of our framework allows for straightforward refinements: expanding the dataset to include a richer variety of architectural designs, as well as incorporating higher‑order geometric moment‑derived MCSIs that also encode local architectural details. These refinements would enable the model to bridge the gap between universally governing nanoscale mechanisms and higher order effects due to architecture-specific nuances, ultimately delivering high-precision predictions across a broad range of nanofibrous foam architectures.
@@ 3.4 Implicit geometric descriptor-based design for independent tuning of mechanical properties
To guide the implicit geometric descriptor-based design of architected VACNT foams, as a first step, we quantified the individual influence of each descriptor on the mechanical properties, through a “quotient of error increase” sensitivity analysis [47] for our ANN model. This method measures how much the network’s error increases when a specific descriptor is excluded, compared to when all descriptors are used. It enables us to assess the importance of each descriptor by evaluating the impact of its absence on the model’s overall performance. For this analysis, we trained the model separately for each mechanical property of interest and computed the MSE and $R^{2}$ values using only one of the two geometric descriptors (Fig. 6(a)). The descriptor importance score $\Omega_{\mathcal{X}}$ is the ratio between the MSE obtained when the descriptor $\mathcal{X}$ is excluded—$\mathrm{MSE}_{\mathcal{X}}$—and the MSE for the original network (Equation 16), providing a measure of the model’s sensitivity to $\mathcal{X}$. Comparing $\Omega$ values across descriptors reveals their relative importance to the ANN model, as shown in Fig. 6(b).
@@
The consistently higher values obtained for $\Omega_{\mathcal{D}}$ across all three specific mechanical properties indicates that $\mathcal{D}$ is the more influential parameter of the two in determining the mechanical properties (Fig. 6(b)). This suggests that, in general, the compactness of the geometries in the architecture should be prioritized to achieve higher specific mechanical properties. However, notably, the influence of $\mathcal{D}$ is lower on $\sigma^{*}$ and even lesser on $E^{*}$, despite $\Omega_{\mathcal{I}_{mc-c}}$ remaining almost unchanged (Fig. 6(b)). This trend is also reflected in the $R^{2}$ values computed for the ANN when one of the geometric descriptors is excluded: the $R^{2}$ values are consistently higher in the absence of $\mathcal{I}_{mc-c}$, although they are reduced for $\sigma^{*}$ and $E^{*}$ (Fig. 6(c)). This finding indicates that $W^{*}$ is predominantly governed by the nanoscale interactions among the nanotubes—captured by $\mathcal{D}$—whereas there is a strong interplay between size effect—captured by $\mathcal{I}_{mc-c}$—and the nanotube interactions in governing the mechanical properties $\sigma^{*}$ and $E^{*}$. Our multi-variate regression analysis with interdependence term included (MVLR-I) further supports this statistically significant interdependence between $\mathcal{D}$ and $\mathcal{I}_{mc-c}$, with p-values of 0.00027 for $\sigma^{*}$ and 0.00015 for $E^{*}$. These findings imply that densely packed and aligned CNTs within confined regions enhance nanotube interactions upon loading, reinforcing the critical role of compactness on $\sigma^{*}$ and $E^{*}$. Therefore, optimizing both $\mathcal{I}_{mc-c}$ and $\mathcal{D}$ in the design of architected VACNT foams could enhance strength and modulus, while designs focused solely on $\mathcal{D}$ may maximize energy absorption. This sensitivity analysis revealed the significance of each descriptor on different mechanical properties, providing a potential foundation for designing VACNT foams to achieve various independent property combinations tailored for a wide range of applications.
@@ 3.5 Generalizable design strategies for target mechanical properties
Understanding how tuning $\tau$ and $\gamma_{i}$ maps onto the mechanical property landscape further refines our strategy. Although $\tau$ may influence $\mathcal{I}_{mc-c}$ differently depending on the region in the $\mathcal{D}$ space, it is intuitive that increasing $\tau$ generally reduces $\sigma^{*}$ and $E^{*}$ owing to the size effect and the reduction in the unit cell population as the unit cell size increases. On the other hand, increasing $\tau$ decreases $\mathcal{D}$ (Fig. 1(d))—the predominant influencer of $W^{*}$—resulting in enhancement of $W^{*}$ (Fig. 6(d)). These trends collectively suggest that by increasing $\tau$, periodic architectures with low $\sigma^{*}$ and $E^{*}$ can be achieved while maintaining high $W^{*}$. Fig. 7(c) and Fig. 7(d) respectively illustrate the $\sigma^{*}$ vs. $W^{*}$ and $E^{*}$ vs. $W^{*}$ property landscapes for cylindrical architectures, highlighting constant $t$ (analogous to $\tau$) contours as predicted by the ANN model. As $t$ increases, the contours shift towards right and stay below the smaller $t$ contours indicating higher $W^{*}$ and lower $\sigma^{*}$ and $E^{*}$. However, along each contour, all three mechanical properties simultaneously increase in the direction where $\mathcal{I}_{mc-c}$ and $\mathcal{D}$ simultaneously decrease due to variations in parameters such as $D_{in}$ and $g$ (analogous to $\gamma_{i}$). Guided by these behaviors, we propose a generalizable, qualitative $\sigma^{*}$(or $E^{*}$) vs. $W^{*}$ property landscape (Fig. 7(e)) for an arbitrary unit cell-based periodic architecture (Fig. 7(e)(Inset)). This landscape presents qualitative constant $\tau$ contours capturing the insights gained from the cylindrical architectures. With the selection of an appropriate $\tau$ to define a target property regime, the $\mathcal{I}_{mc-c}$ and $\mathcal{D}$ can be adjusted by fine-tuning $\gamma_{i}$ to navigate along the contours to realize desired combination of mechanical properties. By leveraging this implicit geometric descriptor enabled-design framework one can simplify the design process and realize systematic pathways for tailoring the mechanical performance of architected VACNT foams across a broad spectrum of applications.
@@ Acknowledgments
MATLAB codes for the generating patterns, computing MCSIs, data processing and ANN training are available on https://github.com/ThevamaranLab/MCSI-enabled-ANN-framefwork
@@
Figure S1 available in Supplementary information (PDF)
Figure S2 available in Supplementary information (PDF)
Figure S3 available in Supplementary information (PDF)
Table S1, S2 , S3, S4 and S5 available in Supplemental information (PDF)